A 2m wide truck is moving with a velocity of 10m/s along a straight road of width 2m. A man starts crossing the road with uniform speed V across the road when the truck is 4m away. The minimum value of V to cross the road safely is ?

Grade:11

1 Answers

Shaswata Biswas
7 years ago
Let the man crosses the road in time t.
Distance to be covered, V.t = 2 m
=> t = 2/V s
Now velocity of the truck is 10 m/s.
So, distance travelled in time = 10.t
Now, for safe crossing, 10t\leq 4 m
\implies 10 \times \frac{2}{V} \leq 4
\implies V \geq 5 m/s
So, the minimum reqiured velocity of the man is 5 m/s.
THANKS